OpenStudy (anonymous):

Multiply: (3x^6y^4)^-1(3x^5y^3)^2

OpenStudy (anonymous):

explanation

OpenStudy (anonymous):

So it looks like this \[(3x^6y^4)^{-1}(3x^5y^3)^2\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes it looks like that I read it wrong the question that is

OpenStudy (anonymous):

\[(3x^6y^4)^{-1}\] is actually \[\frac{ 1 }{ 3x^6y^4 }\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

looks easier now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

great :)

OpenStudy (anonymous):

I got 3 for my answer

OpenStudy (anonymous):

Let's see...

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x^6y^4 } \times (3x^5y^3)^2 = \frac{ (3x^5y^3)^2 }{ 3x^6y^4 }\]

OpenStudy (anonymous):

now you can simplify from there, btw answer is not 3

OpenStudy (anonymous):

\[\frac{ 3^2x ^{10} y^6 }{ 3x^6y^4 }\]

OpenStudy (anonymous):

So you get \[3x^4y^2\]

OpenStudy (anonymous):

i made a mistake at the earlier reply